OpenStudy (gudetamas):

PLEASE HELP WITH FOR MY FINAL. how do i find the inverse of this? f(x)=√x+1

OpenStudy (caozeyuan):

\[f ^{-1}(x)=(x+1)^{2}\]

OpenStudy (caozeyuan):

sorry, (x-1)^2

OpenStudy (caozeyuan):

you take +1 to the other side and square to get rid of the sqrae rroot

OpenStudy (gudetamas):

where do i add +1? to the x?

satellite73 (satellite73):

depends if i tis \[f(x)=\sqrt{x+1}\] or \[f(x)=\sqrt{x}+1\]

OpenStudy (gudetamas):

its the first one

satellite73 (satellite73):

then just like solving say \[\sqrt{x+1}=7\] square, then subtract 1 only this time solve \[\sqrt{y+1}=x\] for \(y\) in the same two steps

OpenStudy (gudetamas):

how did u get 7?

satellite73 (satellite73):

i made it up

satellite73 (satellite73):

can you solve it? can you solve \[\sqrt{x+1}=7\] for \(x\)?

OpenStudy (gudetamas):

you would subtract 1 from 7 right?

satellite73 (satellite73):

no not as a first step you cannot subtract 1 because it is inside the radical, not outside

satellite73 (satellite73):

lets try this we have \[f(x)=\sqrt{x+1}\] what is \(f(48)\)?

OpenStudy (gudetamas):

oh ok so x+1 is 2 right? and then √2 ?

satellite73 (satellite73):

um no, \(x+1=2\) only if \(x=1\) but \(x\) is a variable, it can be any number

satellite73 (satellite73):

for example, \(x\) could be \(48\) what is \(f(48)\)?

OpenStudy (gudetamas):

you want me to square 48?

satellite73 (satellite73):

not for this function no the function is add one, take the square root for example if \[f(x)=\sqrt{x+1}\] then \[f(3)=\sqrt{3+1}=\sqrt4=2\]

satellite73 (satellite73):

and say \[f(8)=\sqrt{8+1}=\sqrt9=3\]

satellite73 (satellite73):

how about \(f(15)\)?

OpenStudy (gudetamas):

f(15) = √15+1 = √16 = 4

satellite73 (satellite73):

yay

satellite73 (satellite73):

how about \(f(48)\)?

OpenStudy (gudetamas):

f(x) = √x+1 = √1 = 1 right?

satellite73 (satellite73):

still asking what is \(f(48)\)

OpenStudy (gudetamas):

f(48) = √48 + 1 = √49 = 7

satellite73 (satellite73):

yes, good so now we know if \(x=48\) then \(f(48)=7\)

satellite73 (satellite73):

now we have to work backwards suppose you have to solve \[\sqrt{x+1}=7\] for \(x\) (we already know the answer is 48) how would we find \(x\) in two steps?

OpenStudy (gudetamas):

subtract 1 from 7?

satellite73 (satellite73):

that give six, that is not going to help finding the 48 is it?

satellite73 (satellite73):

the last step you have was \(\sqrt{49}=7\) above how can you get from 7 to 49?

OpenStudy (gudetamas):

7 times 7?

satellite73 (satellite73):

right, aka \(7^2\)

satellite73 (satellite73):

now how can you get from 49 to 48?

OpenStudy (gudetamas):

subtract by 1

satellite73 (satellite73):

exaclty steps were square, subtract 1 same steps here \[\sqrt{y+1}=x\]solve for \(y\) using the same two steps

OpenStudy (gudetamas):

so you do x^2?

satellite73 (satellite73):

yes that is first, that gives \[y+1=x^2\] next step?

OpenStudy (gudetamas):

and then subtract 1 by x^2

satellite73 (satellite73):

well, i would say "subtract one FROM \(x^2\) but that is probably what you meant how to you write it?

OpenStudy (gudetamas):

yeah sry

satellite73 (satellite73):

ok so what do you get as an answer?\[y=?\]

OpenStudy (gudetamas):

1?

satellite73 (satellite73):

no, not a number, how do you write "square x, subtract 1"

OpenStudy (gudetamas):

ohh ok. x2-1

OpenStudy (gudetamas):

x^2*

satellite73 (satellite73):

yay

satellite73 (satellite73):

that is your answer if \[f(x)=\sqrt{x+1}\] then \[f^{-1}(x)=x^2-1\]

OpenStudy (gudetamas):

thanks so much! i get it now :D

satellite73 (satellite73):

your welcome~! good luck with the next one

OpenStudy (gudetamas):

@satellite73 ccould u help me on this too? f(x) = x^2/2